# VAMP-Net: An Interpretable Multi-Path Framework of Genomic Permutation-Invariant Set Attention and Quality-Aware 1D-CNN for MTB Drug Resistance

### v1

@@ 5.4 Models Experimentation
Table 5: Best hyperparameters of CNN model applied on RIF Drug
@@
The proposed models were evaluated across multiple architectures and hyperparameter configurations. Tables 2, 3, 4, and 5 summarize the best-performing hyperparameters identified through Bayesian optimization. This tuning prioritized a balance between ROC AUC and Accuracy on the validation set, ensuring that models not only fit the training data but also generalized robustly.
Figure 7: Comparison of -Model A- and -Model B- performance without and with variant sequence shuffling for training and validation
In Figure 7, we compare the performance of Model-A and Model-B where Model-B applies data augmentation via variant sequence shuffling. This augmentation fully leverages the SAB block’s capacity, allowing predictions independent of variant position. Training curves confirm that Model-B, likely due to the regularization effect of data augmentation, exhibits greater training stability as evidenced by significantly reduced overfitting and closer convergence between the training and validation curves. On the test set, Model-A achieved the highest performance (ROC AUC: 0.969, Balanced Accuracy: 0.945). While Model-B’s test performance was marginally lower (ROC AUC: 0.964, Balanced Accuracy: 0.941), this minor trade-off in peak performance resulted in a substantially more robust and stable model. This suggests that data augmentation effectively improves generalization and stability with only a negligible decrease in maximum predictive power.
Figure 8: Training (top) and validation (bottom) performance curves of all models with early stopping
Figure 9: ROC AUC Test curve of VAMP-Net vs CNN and MLP
Figure 8 presents the training and validation curves across epochs for eight distinct models (Model-A through Model-H), offering critical insight into their convergence, learning capacity, and generalization ability.
All models exhibit rapid initial learning, achieving 85% to over 90% accuracy within the first 5 to 10 epochs. The models can be broadly categorized into three performance tiers based on their final stabilized validation accuracy and AUC. Models A, B, C, and D demonstrate the highest potential, peaking around 94%. Models G and H show comparatively lower learning capacity for this task, suggesting architectural limitations and insufficient model capacity. Models G and H required the longest training durations, this is a characteristic of models with less efficiency.
Model-A achieved the highest overall training and validation accuracy and AUC within the shortest training time, making it the most efficient high-performer. This demonstrates that Path-2 inclusion significantly improved robustness and that SAB-based transformers outperform classical CNN/MLP representations. The lower performance of Models G and H suggests that these architectures are less suitable for the current task, as shown in Figure 9.
Figure 10: Models test results comparison
The results, visualized in Figure 10, clearly demonstrate the superior performance of the Set Attention Block (SAB)-based transformer architectures. While Model-A achieved the highest overall performance, the marginal variation across other SAB variants underscores the robustness of the core attention mechanism. Notably, the impact of explicit masking appears minimal, as evidenced by the comparable performance of Models C and D, suggesting that the SAB structure inherently handles or ignores padding tokens effectively. In contrast, the significant underperformance of the traditional Multilayer Perceptron (MLP) (Model G) and Convolutional Neural Network (CNN) (Model H) strongly validates the utility of a sequential attention-based representation over standard binary encoding or fixed-filter approaches for this task.
Based on its performance across all evaluation metrics, we selected Model-A as the optimal architecture. To validate the model’s capacity for broad applicability, we applied this model on the RFB drug. On this drug, Model-A demonstrated high efficacy, achieving an AUC of 0.9681 and an Accuracy of 0.9393. Furthermore, the model maintained a well-balanced classification profile, yielding an F1 score of 0.9535, a Recall of 0.9553, and a Precision of 0.9517. These results collectively confirm the model’s excellent generalization capacity across different therapeutic agents.
The comprehensive architectural evaluation leads to the following four key conclusions:
  • Superiority of Attention: Set Attention Block (SAB)-based transformer architectures significantly outperform standard Convolutional Neural Network (CNN) and Multilayer Perceptron (MLP) baselines on presence/absence data.
  • Enhanced Generalization: Augmentation (as employed in Model-B) effectively improves training stability and promotes permutation invariance, leading to more robust models.
  • Architectural Robustness: While empirically minor, the inclusion of attention masks provides critical robustness against padding tokens, formalizing the model’s ability to handle variable-length sequences.
  • Broad Applicability: The high performance achieved on the RFB drug dataset validates the broad applicability and transferability of the SAB-based transformer approach to new, unseen drugs.
5.5 Multi-Path Framework Analysis and Interpretation
@@
This section prensents the results of the Dual-Path Interpretability Framework and leverages the insights to dissect the model’s decision-making process for RIF and RFB resistance prediction. Our analysis demonstrates that the two architectural pathways within VAMP-Net provide distinct yet complementary perspectives on the underlying mechanisms of drug resistance. Specifically, Path-1 (Set Attention Transformer) successfully identified and quantified variant-level importance and epistatic dependencies, thus elucidating the core genetic drivers. In contrast, Path-2 (Quality-Aware CNN) focused on quantifying the influence of technical confidence metrics, revealing how the model performs noise regularization and feature calibration. By integrating the findings from both streams, our framework provides a comprehensive and auditable explanation of the model’s decision process —both biological and technical— that drives the final drug resistance classification.
5.5.1 Set-Transformer Genetic Variant Importance
This section details the quantitative analysis of individual variant contributions to drug resistance, as determined by the Integrated Gradients (IG) method applied to the SAB pathway (Path-1). The IG-derived attribution scores, aggregated across the test cohort, establish an empirically derived ranking of mean importance for all unique ChromPos_Ref>>Alt variants. This analysis serves to validate the model’s ability to prioritize known resistance drivers while simultaneously discovering highly-ranked non-canonical variants that warrant further investigation. The subsequent analyses distinguish between the variant importance profiles learned for Rifampicin (RIF) and Rifabutin (RFB), highlighting the model’s discriminatory power.
A) Drug 1: RIF Resistance Analysis
The analysis of the VAMP-Net pathway for RIF resistance achieved strong validation success, presented in Table 6, demonstrating the framework’s ability to learn and prioritize fundamental genetic mechanisms as reported in the literature (telenti1993detection; ramaswamy1998molecular). The identified variants in rpoB, katG, and other genes align with established resistance mechanisms documented in the WHO catalogue of mutations associated with drug resistance (world2023catalogue).
Table 6: Among 200 Top-Ranked Variants from RIF Resistance Analysis out of 417,635 Unique Variants
Variant Importance Score Gene
@@ 5.5.1 Set-Transformer Genetic Variant Importance
761140_A>>T 0.058 rpoB
761154_TCG>>TTC 0.050 rpoB
761108_GGA>>GTT 0.044 rpoB
760314_G>>T 0.029 rpoB
2155162_ATGCCGC>>ACGCCGT 0.446 katG
2155166_CGC>>CTG 0.043 katG
7581_G>>T 0.050 gyrA
4247574_A>>C 0.036 embB
1849861_G>>T 0.207 ethA
2338202_CAC>>CCAC 0.308 ponA1 (novel)
424791_T>>TAT 0.273 pknB (novel)
The model assigned primary predictive importance to variants within the RNA polymerase $\beta$ subunit (rpoB) gene. Multiple high-confidence variants in Table 6, including 761139_C>>A, 761101_A>>T, and 761140_A>>T, are located within the rpoB gene and fall precisely within the Rifampicin Resistance-Determining Region (RRDR). This alignment with established clinical and molecular biology literature validates the framework’s capacity to learn fundamental resistance pathways (musser1995antimicrobial).
Beyond the primary rpoB target, the VAMP-Net successfully extracted complex, multi-gene resistance signatures. The model assigned high importance scores to variants associated with isoniazid resistance, including mutations in the katG region (2155162_ATGCCGC>>ACGCCGT) and the promoter of ahpC (78519_G>>A) (vilcheze2007mechanism; napier2023large). This demonstrates the framework’s ability to capture complex multi-drug resistance profiles where resistance to one drug often co-occurs with markers for another (jia2021analysis; billows2024large).
A significant finding is the identification of top-ranked variants located in previously uncharacterized genomic regions (Novel loci). The high Importance Scores for these novel variants (such as ponA1 and pknB) suggest the model is uniquely capable of isolating:
  • Rare, High-Effect Mutations: Highly penetrant variants that evade detection by standard frequency-based GWAS methods.
  • Critical Interaction Hubs: Loci that function as epistatic modifiers, significantly altering the expression or effect of primary resistance mutations.
In summary, the high attribution results confirm the robustness and superior discovery potential of the VAMP-Net, which can simultaneously validate known mechanisms and identify high-impact novel loci for targeted functional investigation.
B) Drug 2: RFB Resistance Analysis
The RFB resistance analysis successfully identified the complete spectrum of variants associated with the phenotype (Table 7), confirming the model’s high fidelity in identifying key resistance mechanisms.
Table 7: Among 200 Top-Ranked Variants from RFB Resistance Analysis out of 417,635 Unique Variants
Variant Importance Score Gene
761140_A>>T 6.896 rpoB
761139_C>>A 3.955 rpoB
761140_A>>G 3.321 rpoB
761110_A>>G 3.316 rpoB
761110_A>>T 2.463 rpoB
4247431_G>>T 2.125 embB
7581_G>>A 1.323 gyrA

### v2

@@ 4.3.1 Comparative Performance Against Baseline Architectures
Table 8: Optimized hyperparameter configuration for the CNN baseline. The table outlines the best-performing structural parameters determined via empirical tuning on Rifampicin (RIF) resistance prediction.This specific configuration is referred to as Model H
@@
As illustrated in Figure 12, VAMP-Net consistently establishes a new performance ceiling across all evaluated metrics. The ROC curve analysis reveals that while the CNN achieves a competitive AUC of 0.96, VAMP-Net reaches a superior 0.97, demonstrating a more robust trade-off between sensitivity and specificity. The MLP, lacking specialized spatial or attention-based feature extraction, lags significantly with an AUC of 0.87. The granularity of this superiority is further evidenced by the confusion matrices on the unseen test set (Figure 13). VAMP-Net correctly identifies 92.9% of resistant isolates, a notable improvement over the CNN (89.4%) and the MLP (84.2%). Crucially, as shown in Table 9, the performance gains are statistically validated; VAMP-Net achieved a significant improvement in Accuracy over the CNN ($1.35\%$, $p<0.001$) and the MLP ($8.48\%$, $p<0.001$). The 95% Confidence Intervals for VAMP-Net remain narrow (e.g., Accuracy: $[0.9383,0.9554]$), underscoring the model’s high reliability and precision in a clinical diagnostic context.
Figure 12: Comparative performance of VAMP-Net against baseline architectures on the test set. (Left) Receiver Operating Characteristic (ROC) curves demonstrating the trade-off between true positive and false positive rates. (Right) Detailed test set metrics across six performance indicators.
Figure 13: Comparative confusion matrix analysis on the unseen test set. This figure provides a head-to-head performance breakdown between the proposed VAMP-Net architecture and two baseline models: a Convolutional Neural Network (CNN) and a Multi-Layer Perceptron (MLP).
Table 9: VAMP-Net overall performance and statistical significance comparison against baselines.
Metric Value 95% CI p-value (vs Baselines)
Accuracy 0.9474 [0.9383, 0.9554] $0.001^{**}$
F1-score 0.9555 [0.9476, 0.9627] $0.001^{**}$
ROC-AUC 0.9691 [0.9612, 0.9762] 0.074
∗∗Significant at $p<0.001$ relative to CNN and MLP baselines.
4.3.2 Comparative Evaluation of Fusion Paradigms
The architectural efficacy of VAMP-Net’s late-stage adaptive fusion was benchmarked against a traditional Early Fusion baseline, applied to RIF drug, to quantify the impact of integration timing on model performance. As summarized in Table 10, VAMP-Net significantly outperformed the Early Fusion paradigm across all evaluated metrics. Most notably, VAMP-Net achieved transformative gains in Specificity (+48.8%), Balanced Accuracy (+32.6%), and ROC-AUC (+30.9%) compared to the early-integration approach. The Early Fusion model exhibited a profound performance asymmetry, demonstrating high Sensitivity (0.805) but catastrophic Specificity (0.425). This imbalance suggests a systematic failure to distinguish true genomic resistance markers from stochastic noise. In contrast, VAMP-Net maintained a harmonized performance profile, achieving a Recall of 0.968 and a Specificity of 0.913.
The empirical results indicate that forcing inter-modality interaction at a premature stage—prior to dedicated contextual modeling—is suboptimal for complex genomic tasks. In the Early Fusion setting, token-level embeddings and structured features are concatenated and projected before entering the Set Attention Blocks (SABs). Our findings suggest that this approach leads to the propagation of unrefined signal noise, which the subsequent attention mechanisms struggle to resolve, as evidenced by the severe specificity deficit. VAMP-Net’s superiority is rooted in its late-stage adaptive fusion philosophy. By postponing integration until after modality-specific encoding, each pathway is permitted to develop robust internal representations. This allows the SAB encoder to model inter-variant dependencies within a more stable feature space.
Table 10: Comparative analysis of fusion architectural paradigms. This table benchmarks the proposed VAMP-Net multi-path integration (late-stage adaptive fusion) against a traditional Early Fusion baseline.
Metric Early Fusion VAMP-Net Fusion
Accuracy 0.649 0.946
F1-score 0.729 0.955
Sensitivity (Recall) 0.805 0.968
Specificity 0.425 0.913
Precision (PPV) 0.667 0.941
NPV 0.605 0.953
Balanced Accuracy 0.615 0.941
ROC-AUC 0.655 0.964
4.3.3 Comparative SOTA
To provide a more faithful comparison, we separate prior work by paper and retain only the metrics explicitly reported in each study. For every drug shared with our benchmark, we compare the published result against the best-performing VAMP-Net seed for that drug on our validation protocol. Because these studies use different cohorts, split strategies, and prediction formulations, the following tables should be interpreted as paper-specific contextual comparisons rather than direct leaderboard claims.
Table 11: Paper-specific comparison with the XGBoost study in Diagnostics [31]. The original paper reports Sensitivity, Specificity, Precision, F1-score, and Accuracy for EMB, INH, and RIF.
Drug Our Seed Sens. (Paper) Sens. (Ours) Spec. (Paper) Spec. (Ours) Prec. (Paper) Prec. (Ours) F1 (Paper) F1 (Ours) Acc. (Paper) Acc. (Ours)
EMB 42 0.97 0.8166 0.97 0.8885 0.89 0.9673 0.93 0.8856 0.97 0.8309
INH 42 0.90 0.9382 0.99 0.9439 0.97 0.9441 0.94 0.9411 0.97 0.9410
RIF 2024 0.94 0.9360 0.96 0.9501 0.90 0.9640 0.92 0.9498 0.96 0.9418
Compared with the XGBoost results reported in Diagnostics, VAMP-Net shows a more favorable precision profile across all shared drugs, with particularly strong gains for EMB and RIF, indicating that the proposed architecture is more conservative in assigning resistant calls while preserving competitive sensitivity. The main trade-off appears for EMB, where the prior study reports higher sensitivity and accuracy, suggesting that EMB remains the most challenging phenotype in our setting and may be more sensitive to cohort composition and decision-threshold effects.
Table 12: Paper-specific comparison with the MIC-prediction machine learning system reported in PLOS Computational Biology [10]. Only metrics shared with our evaluation are shown; the paper reports PPV, which is reported here as the published precision-equivalent quantity.
Drug Our Seed AUC (Paper) AUC (Ours) Sens. (Paper) Sens. (Ours) Spec. (Paper) Spec. (Ours) PPV/Prec. (Paper) Prec. (Ours) Acc. (Paper) Acc. (Ours)
EMB 42 0.965 0.9236 0.931 0.8166 0.894 0.8885 0.704 0.9673 0.902 0.8309
INH 42 0.983 0.9696 0.950 0.9382 0.981 0.9439 0.980 0.9441 0.966 0.9410
RIF 2024 0.992 0.9712 0.971 0.9360 0.973 0.9501 0.962 0.9640 0.972 0.9418
RFB 2024 0.988 0.9733 0.955 0.9782 0.957 0.8882 0.918 0.8874 0.956 0.9309
Relative to the MIC-oriented model from PLOS Computational Biology, VAMP-Net remains competitive but reveals a mixed pattern. Our model achieves its closest agreement on INH and RFB, while the largest performance gaps are observed for EMB and RIF, especially in AUC and accuracy. This suggests that VAMP-Net is strongest when technical-confidence recalibration contributes substantially to the final decision, but that MIC-prediction frameworks trained under different targets and evaluation designs can retain an advantage on some first-line drugs.
Table 13: Paper-specific comparison with LLMTB, the BERT-based large language model reported in Bioinformatics [45]. The paper reports positive-class F1-scores on the shared drugs.
Drug Our Seed F1 (LLMTB) F1 (Ours)
EMB 42 0.8273 0.8856
INH 42 0.9361 0.9411
RIF 2024 0.9094 0.9498
RFB 2024 0.8913 0.9306
Against LLMTB, VAMP-Net improves the reported F1-score for all four shared drugs, with the clearest gains on EMB, RIF, and RFB. This pattern supports the value of combining permutation-invariant biological modeling with an explicit quality-aware branch, rather than relying solely on sequence-style language representations. The smaller margin on INH further suggests that both architectures are already highly effective for this phenotype, leaving less headroom for separation.
4.4 Path-2 Contribution and Case Studies:
To better understand how the quality-aware Path-2 branch influences final predictions, we inspected representative test cases in which the fused model either corrected an overconfident Path-1 decision or amplified a borderline but correct Path-1 prediction. Table 14 summarizes these case studies using the resistance probability predicted by the variant-only pathway (Path-1) and by the full fused model. We additionally report the probability shift $\Delta P=P_{\mathrm{Full}}(\mathrm{Resistant})-P_{\mathrm{Path\mbox{-}1}}(\mathrm{Resistant})$, together with summary statistics of the learned fusion gate. These examples show that the VCF-aware Path-2 branch is not merely an auxiliary source of features, but an active recalibration mechanism for Path-1’s variant-based decision. In the corrected cases, Path-1 predicted resistance with very high confidence ($P(R)\approx 0.91$), yet the fused model reduced this probability to 0.040.040.060.06, corresponding to a large decision shift of roughly 0.86-0.86. This is not a marginal adjustment; it is a strong reversal of an initially overconfident prediction.
Table 14: Ablation case studies of VCF-aware Path-2 contribution. All samples feature a variant count of 743. “Corrected” denotes cases where Path-B reversed an incorrect Path-1 prediction, while “Amplified” denotes Path-2 reinforcing a correct class by shifting borderline predictions away from the decision boundary.
True Label Path-1 $P(R)$ Full $P(R)$ $\Delta P$ Effect Principal Explanation
Susceptible 0.9190 0.0521 0.867-0.867 Corrected High FRS but low genotype-confidence and depth support made Path-1’s resistance call unreliable.
Susceptible 0.9049 0.0438 0.861-0.861 Corrected Path-A overcalled resistance; Path-2 utilized VCF-support signals to downweight the prediction.
Susceptible 0.9185 0.0592 0.859-0.859 Corrected Low-confidence and low-depth evidence triggered the fused model to reverse the initial resistant call.
Susceptible 0.4579 0.0774 0.381-0.381 Amplified Borderline sample pushed further into the susceptible class, increasing overall model confidence.
Resistant 0.5491 0.9028 +0.354+0.354 Amplified Combined variant identity with consistent support features to strengthen a true resistant call.
Resistant 0.5495 0.8826 +0.333+0.333 Amplified Reinforced the resistant signal, moving a borderline Path-1 output safely away from the decision boundary.
Inspection of the associated feature profiles clarifies why this occurs. In these corrected cases, FRS remained high, indicating that the called alleles were internally supported by reads, but the broader VCF context, particularly GT_CONF_PERCENTILE, depth-related features, and filtered depth support, remained weak. Thus, although Path-1 detected a mutation pattern that appeared resistance-like, Path-2 identified insufficient technical evidence to trust that pattern as a robust resistance signal. In this sense, Path-2 behaves as a reliability-aware filter: variant presence alone is not treated as sufficient, and the model instead asks whether the surrounding sequencing evidence is strong and consistent enough to justify a resistant call. The amplified cases demonstrate the complementary behavior. When Path-1 produced borderline predictions, Path-2 was able to push the fused output farther toward the correct class. For truly resistant isolates, the full model increased the resistance probability from about $0.55$ to 0.880.880.900.90, showing that the quality-aware branch does not only suppress unreliable signals, but can also strengthen biologically meaningful signals when the VCF support is coherent. Overall, these case studies show that the fused model can correct Path-1 false positives by incorporating genotype confidence, depth, and coverage-derived evidence, thereby reducing spurious resistance probabilities by as much as $0.87$ and yielding more reliable susceptibility calls. Figure 14 provides a visual quantification of the "corrective" and "amplificatory" influence of the Path-2 quality-aware branch. By plotting the resistance probability of Path-1 against the final model output, we can observe the specific instances where sequencing metadata fundamentally altered the diagnostic outcome.
Figure 14: Probabilistic recalibration of resistance calls through multi-path fusion. The scatter plot illustrates the shift in resistance probability $P(\text{Resistant})$ when transitioning from the variant-only model (Path-1) to the fully integrated VAMP-Net architecture (Full Model). Points in the lower-right quadrant represent "Corrected" cases, where Path-1 initially assigned high resistance probabilities (>0.5>0.5) that were subsequently reversed by Path-2’s quality-aware branch to confident susceptible calls (<0.2<0.2). Conversely, points in the upper-left quadrant denote "Amplified" cases, where Path-2 reinforced borderline or incorrect susceptible predictions into high-confidence resistance calls. The significant deviation from the "No change" diagonal underscores Path2’s role as a critical reliability filter that identifies and corrects overconfident predictions based on underlying sequencing evidence.
4.5 Multi-Path Network Interpretation and Epistasis
@@
This section prensents the results of the Dual-Path Interpretability Network and leverages the insights to dissect the model’s decision-making process for RIF, RFB, INH and EMB resistance prediction. Our analysis demonstrates that the two architectural pathways within VAMP-Net provide distinct yet complementary perspectives on the underlying mechanisms of drug resistance. Specifically, Path-1 (Set Attention Transformer) successfully identified and quantified variant-level importance and epistatic dependencies, thus elucidating the core genetic drivers. In contrast, Path-2 (Quality-Aware CNN) focused on quantifying the influence of technical confidence metrics, revealing how the model performs noise regularization and feature calibration. By integrating the findings from both streams, our network provides a comprehensive and auditable explanation of the model’s decision process, both biological and technical, that drives the final drug resistance classification.
4.5.1 Mechanistic Interpretation and Variant Attribution
To quantify the individual genetic contributions to drug resistance, we applied the Integrated Gradients (IG) method to the SAB pathway. As detailed in Table 15, the resulting importance scores establish a hierarchical signature that validates the network’s ability to autonomously rediscover established molecular pathways while proposing high-impact novel candidates.
Table 15: Top-ranked genomic variants and associated genes identified by VAMP-Net. The table lists high-importance variants prioritized by the model’s attention mechanism for RIF, RFB, INH and EMB. While the model successfully recovers established resistance determinants in the rpoB, katG, and ethA genes, it also identifies several "novel" high-impact loci in genes such as ponA1, pknB, and mmpL5.
Drug Variant Importance Score Gene
RIF 2155162_ATGCCGC>>ACGCCGT 0.446 katG
2338202_CAC>>CCAC 0.308 ponA1 (novel)
424791_T>>TAT 0.273 pknB (novel)
1849861_G>>T 0.207 ethA
@@ 4.5.1 Mechanistic Interpretation and Variant Attribution
RFB 761140_A>>T 6.896 rpoB